OpenStudy (anonymous):

if f(x)=2x-4 & g(x)=2^2 what is f(g(x)) ?

OpenStudy (anonymous):

A) x^2-4 B) x^2+2x-4 C) 2x^3-4 D)2x^2-4

OpenStudy (anonymous):

Is g(x) = 2^2 or x^2?

OpenStudy (anonymous):

x^2 sorry

OpenStudy (anonymous):

Alright. The answer will then be: F(x) = 2x - 4 G(x) = x^2 F(G(x)) = F(x^2) = 2x^2 - 4 Which is D.

OpenStudy (anonymous):

thank you(: